If you don't mind helping me out with my homework: I would be very grateful!
x² + bx + c = (x + p)(x + q), where pq = b and p + q = c.

True

False

Answer by ikleyn(52848)

.
If you don't mind helping me out with my homework: I would be very grateful!
x² + bx + c = (x + p)(x + q), where pq = b and p + q = c.
True
False
-----------------------------------------------------------

False.

The correct identities are pq = c and p + q = b.